OpenStudy (anonymous):

The sum of four consecutive integers is 74. What is the first integer? 16 17 18 19

OpenStudy (abdullahm):

Do you know what consecutive integers mean?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Not consecutive.

OpenStudy (abdullahm):

"Consecutive integers are integers that follow each other in order. They have a difference of 1 between every two numbers. In a set of consecutive integers, the mean and the median are equal. If n is an integer, then n, n+1, and n+2 would be consecutive integers. " ~ https://simple.wikipedia.org/wiki/Consecutive_integer

OpenStudy (anonymous):

okay.

OpenStudy (abdullahm):

So, we can make an equation since we know they are consecutive integers... Also by seeing the example in the last paragraph I copied and pasted above. So we have: \(\sf\Large x + (x+1) + (x+2) + (x+3) = 74\)

OpenStudy (abdullahm):

Simplify that and solve for x.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x+x+1+x+2+x+3=74 (x+x+x+x)+(1+2+3)=74 4 x's add 1+2+3=6 4x+6=74 subract 6 from both sides 4x=68 Divide both sides by 4 4x/4=68/4 And my answer was x=17

OpenStudy (abdullahm):

Correct, good job! :D

OpenStudy (anonymous):

:)